uty lat thè unit has exactly four rooms, given that it has at le two rooms. The conditional probability that the unit has at how to calculate d.
Answer #1

Solution:-

Possible outcomes=HHH, HTH, THH, TTH, HHT, HTT, THT, TTT

event the first toss is heads=HHH, HTH, HHT, HTT

P(A)=P(event the first toss is heads)=4/8

event that the total number of heads is 1=HTT, THT, TTH

P(C)=P(event that the total number of heads is 1)=3/8

Now,

(event the first toss is heads) intersection (event that the total number of heads is 1)=HTT,

P(A ∩ C)=1/8

so, P(C/A)=P(P(A ∩ C)/P(A)

=(1/8)/(4/8)

=0.25